HIGH-RESOLUTION INFRARED SPECTRUM OF THE 3ν_{9} −ν_{9} HOT BAND OF HNO_{3}: STUDY OF CORI-OLIS COUPLED 3ν_{9} AND THE NEARBY DARK STATE

\underline{W. F. WANG}, J. MARCOS SIROTA, Joint Center for Earth Systems Technology, University of Maryland Baltimore County, 1000 Hilltop Circle, Baltimore, Maryland 21250, USA; P. P. ONG, and E. C. LOOI, Department of Physics, Faculty of Science, National University of Singapore, Lower Kent Ridge Road, Singapore 119260, Singapore.

The high-resolution infrared spectrum of HNO_{3} in the region of 805-853 cm^{−1} has been measured and assigned for the 3ν_{9} −ν_{9} hot band. The significant splittings in the upper state 3ν_{9} arising from the torsional motion of H around ON were observed and identified. It was found that 3ν_{9}, which is around 1289 cm^{−1}, was involved in a strong Coriolis resonance with a dark state near 1301 cm^{−1}, apart from the slightly higher vibrational state ν_{4}. To account for the corresponding Coriolis perturbations as well as the torsional splittings observed in the 3ν_{9}−ν_{9} band, a comprehensive line-position fit was performed by simultaneously adjusting three sets of constants for 3ν_{9,1}, 3ν_{9,2} and the dark state. Accurate rovibrational constants were obtained giving a fit of 947 unperturbed and 146 perturbed lines to the rms accuracy of 0.00062 cm^{−1} and 0.00175 cm^{−1}, respectively. The interacting dark state, found at 1300.7331(19) cm^{−1}, can be assigned to 2ν_{6} which proves to be responsible for the previously unjustified ^{a} ΔK_{c} = 6 resonance observed in the strong ν_{3} band.

aA. Perrin, O. Lado-Bordowsky, and A. Valentin, Mol. Phys. 67, 249-270, 1989